OpenStudy (anonymous):

the question to be solved should be written 3xy=4 + 4x. The y isn't in brakets

OpenStudy (anonymous):

3XY = 4 + 4X Y = (4+4X)/3X Y= 4/3X + 4/3

OpenStudy (anonymous):

Well, what answer are you expecting? What is it that is to be solved?